Investment Center Net Income Average Assets
Cameras and camcorders $ 6,450,000 $ 29,000,000
Phones and communications 2,982,000 21,300,000
Computers and accessories 1,050,000 17,400,000

Assume a target income of 14% of average invested assets.

Required:
Compute residual income for each division. (Enter losses with a minus sign.)
Answer #1

residual income=Net income-(Average Assets*Target rate of return)

Residual income
Cameras and camcorders 6,450,000-(29,000,000*14%)=$2,390,000
Phones and communications 2,982,000-(21,300,000*14%)=0
Computers and accessories 1,050,000-(17,400,000*14%)=($1,386,000)(Negative).